OpenStudy (anonymous):

For what value of k are the lines 2x+3y=4k and x-2ky=7 parallel? I really have no idea how to deal with the extra variable.

Directrix (directrix):

Something to remember: Parallel lines have the same slope.

Directrix (directrix):

2x + 3y = 4k 3y = -2x + 4k y = (-2/3)x + (4/3)k Slope is -2/3

Directrix (directrix):

Will you find the slope of this line? x-2ky=7 @CasidyCat

OpenStudy (anonymous):

@Directrix That was the problem. I couldn't find the slope for the second equation.

Directrix (directrix):

x - 2ky = 7 2ky = x - 7 y = (1/(2k))x - (7/(2k)) Slope is 1/(2k)

Directrix (directrix):

If the lines are parallel, the slopes are equal: 1/(2k) = -2/3 2k = -3/2 4k = -3 k = -3/4 k = -0.75

Directrix (directrix):

Input from @jeff1949

Directrix (directrix):

@CasidyCat Any questions?

OpenStudy (anonymous):

Nope, thank you! I was just getting confused as to what I should consider the k, but I understand now.

Directrix (directrix):

Alrighty, then.